# PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models

### v1

@@ 3.4 Rationales Decoder Module
    Exponential Mechanism Decoder Strategy. In the first decoding strategy, which corresponds to Exponential Mechanism Encoder strategy. Here, we utilize In-Context Learning(ICL) Dong et al. (2022) Tong et al. (2023) with the SLM to decode the perturbed rationales. we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the SLM to prompt the generation of rationales, where $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt and $r^{p}$ represents perturbed rationales generated from LLM. $(p^{p},r^{p})_{i}$ can be viewed as an example for SLM in ICL. This allows the SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
@@
    Encoder-Decoder Decoder Strategy. In the second decoding strategy, which corresponds to Encoder-Decoder Encoder strategy. The rationales decoder module also use the same the Encoder-Decoder SLM with Section 3.2.
@@
    The Decoder objective can be formulated as follows:
@@
    where $\mathcal{L}_{\text{Decoder}}$ is the rational decoder loss, and $\ell_{\text{CE}}$ is the cross-entropy loss.
@@
    We approach the training of the Encoder-Decoder SLM as a multi-task learning problem encompassing both the encoding and decoding training processes. The multi-task learning objective can be formulated as follows:
    $\mathcal{L}_{1}=\alpha\mathcal{L}_{\text{Encoder}}+(1-\alpha)\mathcal{L}_{\text{Decoder}}$ (5)
    where $\alpha$ is the hyperparameters that control the weight of encoder and decoder loss.
As illustrated in Figure 2, we can observe an exemplary comparison between the perturbed rationales from LLM and its decoded rationales from SLM in Step 3 and Step 4. It’s worth noting that although the SLM has the ability to generate aligned rationales independently, the quality often falls short due to its limited capabilities. By leveraging the perturbed rationales, we effectively transfer the powerful capabilities of the server-side LLM to enhance the Encoder-Decoder SLM, thereby improving the overall quality of the generated rationales.
Figure 4: Rationales Decoder Module.
Algorithm 1 PDSS
1:
2:$T$: total number of rounds;
3:$\mathcal{P}$: encoding training datasets;
$\mathcal{R}$: decoding training datasets;
5:$\mathcal{D}$: task-Spec training datasets;
6:$\eta_{\phi}$: learning rate of Encoder-Decoder SLM;
7:$\eta_{\omega}$: learning rate of Task-Specific SLM.
8:$g_{\phi}$, $f_{\omega}$.
9$\triangleright$ Multi-Task Training for Encoder-Decoder SLM based on Public Datasets $\mathcal{P}$ and $\mathcal{R}$.
10:for each epoch $t\in[T]$ do
11:     $\phi^{t+1}\leftarrow\phi^{t}-\eta_{\phi}\bigtriangledown\mathcal{L}_{1}$.
12:end for
13$\triangleright$ Generated $p^{p}$ using the updated Encoder.
14:$p^{p}=SLM_{Encoder}(p)$.
15$\triangleright$ Generated perturbed rationales from LLM on the server.
16:$r^{p}=LLM(p^{p})$.
17$\triangleright$ Decoded perturbed rationales using the updated Encoder-Decoder SLM.
18:$r=SLM_{Decoder}(r^{p})$.
19$\triangleright$ Multi-Task Training for Task-Specific SLM based on Datasets $\mathcal{D}$.
20:for each epoch $t\in[T]$ do
21:     $\omega^{t+1}\leftarrow\omega^{t}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$.
22:end for
3.5 Task-Specific SLM Training Module
@@
We conceptualize learning with rationales as a multi-task learning problem, as illustrated in Figure 5. Specifically, we train the model $f_{\omega}(x_{i})\to(y_{i},r_{i})$ to accomplish not just the prediction of task labels but also the generation of the corresponding rationales based on textual inputs. This multi-task training ensures that our model not only produces accurate predictions but also provides insightful justifications for its decisions. By doing so, we enhance the transparency and explainability of the model. The multi-task learning objective can be formulated as follows:
$\mathcal{L}_{2}=\beta\mathcal{L}_{\text{Label}}+(1-\beta)\mathcal{L}_{\text{Rationale}}$ (6)
@@
where $\ell_{\text{CE}}$ is the cross-entropy loss, $f_{\omega}(.)$ is the Task-Specific SLM model, and $\beta$ is the hyperparameters that control the weight of label prediction loss and rationale generation loss.
Figure 5: Task-Specific SLM Training Module.
@@ 4.1 Setup
We have established a scenario to evaluate the performance of the PDSS framework across a range of text generation tasks. This setup involves a client-server architecture, where the client holds two downstream SLMs :an Encoder-Decoder SLM, which specializes in encoder-decoder functionalities and a Task-Specific SLM, tailored for specific tasks. On the server-side, we host a LLM for more general and powerful text generation capabilities. Specifically, we have chosen Qwen-14BBai et al. (2023) as LLM, while both SLMs are Qwen-0.5BBai et al. (2023). Table 1 outlines the detailed configurations of both the LLM and the SLMs.
Setting Server Client Client
Model Type LLM Encoder-Decoder SLM Task-Specific SLM
Model Name Qwen-14B Qwen-0.5B Qwen-0.5B
Parameters(Billion) 14 0.5 0.5
Table 1: LLM and SLMs Setting of PDSS.
Datasets and Evaluation Metrics. We conduct a comprehensive evaluation of PDSS on 4 QA datasets. Specifically, we include CommonsenseQA(CQA) Talmor et al. (2018), OpenBookQA(OBQA) Mihaylov et al. (2018), BoolQ Clark et al. (2019), ARC-EClark et al. (2018). For these datasets, we primarily use Accuracy as the evaluation metric.
Baselines. Since we incorporate two distinct strategies in the prompt encoder module and rationales decoder module, we denote PDSS method with the Exponential Mechanism Strategy as PDSS-EM and PDSS method with the Encoder-Decoder Strategy as PDSS-ED. We conduct a comparative analysis to evaluate the performance of our PDSS framework, which comprises both PDSS-EM and PDSS-ED.
@@
  • Standalone, where the client independently fine-tunes its local model using its own private dataset;
  • DSSHsieh et al. (2023), where the client fine-tunes its local model by distilling step-by-step LLM method without privacy-preserving.
4.2 Overall Performance Evaluation
In this section, we undertake a comprehensive analysis of the task performance of PDSS. We assess both the PDSS-EM and PDSS-ED methods against other baselines on Task-Specific SLM across various privacy budgets, denoted by $\epsilon$.
The results, as presented in Table 2, clearly illustrate that both PDSS-EM and PDSS-ED exhibit significantly better performance when compared to FewShot-SLM and Standalone methods. With an increase in the privacy budget $\epsilon$, both the performance of PDSS-EM and PDSS-ED have risen notably. Furthermore, PDSS-ED demonstrates notably superior performance compared to PDSS-EM under the same privacy budget $\epsilon$ Specifically, under a privacy budget of $\epsilon=3$, PDSS-EM surpasses the Standalone method by 3.4% and 17% in the CQA and OBQA datasets, respectively, while PDSS-ED outperforms it by 5.2% and 22.4%. Similarly, when the privacy budget is increased to $\epsilon=10$, PDSS-EM exceeds the Standalone approach by 6.3% and 21.6% within the CQA and OBQA datasets, respectively, and PDSS-ED beats it by 7.2% and 28.6%. Remarkably, across all datasets evaluated, when the privacy budget is set to $\epsilon=10$, PDSS achieves comparable performance to DSS, highlighting its efficacy and versatility in balancing privacy and utility.
Method CQA OBQA BoolQ ARC-E
@@ 4.2 Overall Performance Evaluation
DSS 59.3 55.1 80.5 57.6
PDSS-EM($\epsilon=1$) 57.7 49.2 80.1 52.3
PDSS-EM($\epsilon=3$) 57.6 50.8 79 52.6
PDSS-EM($\epsilon=5$) 58.8 53.2 80 55.3
PDSS-EM($\epsilon=10$) 59.2 52.8 80.2 56.2
PDSS-ED($\epsilon=1$) 58.2 50.8 80.3 56.4
PDSS-ED($\epsilon=3$) 58.6 53.1 80.2 56.5
PDSS-ED($\epsilon=5$) 58.3 53.4 80.4 56.3
PDSS-ED($\epsilon=10$) 59.7 55.8 80.7 57.9
@@
Table 2: We compare the performance of Task-Specific SLM trained with PDSS-EM and PDSS-ED across different privacy budgets $\epsilon$ against the Task-Specific SLM trained using baseline methods.
4.3 Reducing Training Data Evaluation
In this section, we conduct an in-depth analysis to explore the influence of training data size on model performance. We compare the PDSS method with the Standalone approach, varying the amount of training data used. Table 3 provides a clear illustration of how PDSS(with $\epsilon=3$) consistently outperforms the Standalone method.
Remarkably, PDSS achieves superior performance even with significantly fewer training samples compared to Standalone. More specifically, when trained on merely 75% of the complete CQA, OBQA, and BoolQ datasets, both PDSS-EM and PDSS-ED surpasses the performance of Standalone fine-tuning that has been trained on the entirety of these datasets. Likewise, by using only 50% of the full ARC-E dataset, PDSS-EM exceeds the results achieved by Standalone fine-tuning on the complete dataset. Furthermore, PDSS-ED exhibits significantly better performance than PDSS-EM across various dataset sizes (ranging from 25% to 100%). The results indicate that PDSS is capable of extracting more valuable information from smaller datasets, making it a promising approach in data-scarce environments.
Task Method 25% 50% 75% 100%
CQA PDSS-EM 49 53.5 56.7 57.6
PDSS-ED 54.2 54.6 56.1 58.6
Standalone - - - 55.7
OBQA PDSS-EM 34.8 42.2 45.6 50.8
PDSS-ED 41.4 43.6 50.6 53.1
Standalone - - - 44.2
BoolQ PDSS-EM 63 74 78.7 79
PDSS-ED 72.8 77.6 79.1 80.2
Standalone - - - 78.4
ARC-E PDSS-EM 45.3 52.2 53.1 53.8
PDSS-ED 48 49.7 55.9 56.5
Standalone - - - 50.3
Table 3: We compare the performance of Task-Specific SLM trained with PDSS-EM($\epsilon=3$) and PDSS-ED($\epsilon=3$) against Standalone, across a range of dataset sizes from 25% to 100%. The ’-’ indicates a method does not apply to the corresponding dataset sizes.
4.4 Perturbed Rationales Evaluation
@@
In this section, we focus on analyzing the quality of the perturbed rationales($r^{p}$) generated from the perturbed prompt of LLM based on PDSS-EM and PDSS-ED methods and compare them with the rationales($r$) generated from raw prompt of the LLM. To evaluate the similarity between $r^{p}$ and $r$, we use TokenRatio metric. A higher TokenRatio indicates a greater degree of similarity between the perturbed and original rationales. For more details about TokenRatio, please refer to Appendix C.
@@
As shown in Table 4, with an increase in the privacy budget $\epsilon$ and a corresponding decrease in perturbation, both the TokenRatio of PDSS-EM and PDSS-ED have risen notably. Furthermore, in most of tasks, the TokenRatio of PDSS-ED is higher than that of PDSS-EM in the same level of privacy budget $\epsilon$. The experimental results confirm that the TokenRatio observed in the perturbed rationales produced by both PDSS-EM and PDSS-ED, positively correlate with the privacy budget $\epsilon$. This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the perturbed rationales become more similar to the original rationales. This finding is significant as it demonstrates the trade-off between privacy protection and the utility of the generated rationales.
Method CQA OBQA BoolQ ARC-E
PDSS-EM($\epsilon=1$) 19.8 26.2 26.6 24.6
PDSS-EM($\epsilon=3$) 29.2 37.2 35.5 33.9
PDSS-EM($\epsilon=5$) 48.8 59.6 55.2 53.9
PDSS-EM($\epsilon=10$) 69.7 72 74.6 68.2
PDSS-ED($\epsilon=1$) 26.7 33.1 29.7 31
PDSS-ED($\epsilon=3$) 33.1 40.9 40.4 42.9
PDSS-ED($\epsilon=5$) 49.6 61 57.5 63.5
PDSS-ED($\epsilon=10$) 57.2 68.3 68 74.2
@@
Table 4: We conduct a comparative analysis to assess the perturbed rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.
4.5 Decoded Rationales Evaluation
@@
In this section, we delve into the quality analysis of the decoded rationales produced by the rationales decoder module based on PDSS-EM and PDSS-ED methods. We compare these decoded rationales against those generated directly from raw prompt of the LLM. We utilize the TokenRatio metric to assess their similarities.
@@
As shown in Table 5, in contrast to FewShot-SLM, it becomes apparent that the decoded rationales’ quality based on PDSS-EM and PDSS-ED methods isn’t solely reliant on the locally decoded SLM. The perturbed rationales crafted by the LLM indeed fulfill their intended purpose. When juxtaposed with Table 4, it’s clear that at comparable $\epsilon$ levels, the TokenRatio for the decoded rationales surpass those of the perturbed rationales in the PDSS-EM and PDSS-ED methods. This underscores the effectiveness of the rationales decoder module in the PDSS-EM and PDSS-ED methods. Furthermore, with the increase of the privacy budget $\epsilon$, the TokenRatio for the decoded rationales generated by both PDSS-EM and PDSS-ED have increased significantly. This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the decoded rationales become more similar to the original rationales. For more details about comparative analysis of perturbed rationales and decoded rationales, please refer to Appendix D.
Method CQA OBQA BoolQ ARC-E
FewShot-SLM 43.3 43.4 51.9 42.6
PDSS-EM($\epsilon=1$) 38.3 37.1 38.4 41.5
PDSS-EM($\epsilon=3$) 41.9 41.3 41.7 45.6
PDSS-EM($\epsilon=5$) 53.1 54 55 58.3
PDSS-EM($\epsilon=10$) 71.1 63 73.6 70.4
PDSS-ED($\epsilon=1$) 57.2 53.4 45.2 57.5
PDSS-ED($\epsilon=3$) 59 55.1 48 59.4
PDSS-ED($\epsilon=5$) 59.8 59.5 55.7 65.5
PDSS-ED($\epsilon=10$) 62 62.3 63.4 70.1
@@
Table 5: We conduct a comparative analysis to assess the decoded rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.

### v2

@@ 3.3 Perturbed Rationales Decoder
    Exponential Mechanism Decoder Strategy. In the first decoding strategy, which corresponds to Exponential Mechanism Encoder strategy. Here, we utilize In-Context Learning (ICL)  dong2024survey; tong2025inferdpt with the Encoder-Decoder SLM to decode the perturbed rationales. we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the Encoder-Decoder SLM to prompt the generation of rationales, where $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt and $r^{p}$ represents perturbed rationales generated from LLM. $(p^{p},r^{p})_{i}$ can be viewed as an example for Encoder-Decoder SLM in ICL. This allows the Encoder-Decoder SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
@@
    Adaptive Exponential Mechanism Decoder Strategy. In the second decoding strategy, which corresponds to Adaptive Exponential Mechanism Encoder strategy. The rationales decoder module also use the same the Encoder-Decoder SLM with Section 3.1.
@@
    The Rationales Decoder objective can be formulated as follows:
@@
    We approach the training of the Encoder-Decoder SLM as a multi-task learning problem encompassing both the encoding and decoding training processes.
    The multi-task learning objective for the Encoder-Decoder SLM can be formulated as follows:
    $\mathcal{L}_{1}=\mathcal{L}_{\text{Encoder}}+\mathcal{L}_{\text{Decoder}}$ (6)
As illustrated in Figure1(b), we can observe an exemplary comparison between the perturbed rationales from LLM and its decoded rationales from SLM in Step 3 and Step 4. It’s worth noting that although the SLM has the ability to generate aligned rationales independently, the quality often falls short due to its limited capabilities. By leveraging the perturbed rationales, we effectively transfer the powerful capabilities of the server-side LLM to enhance the Encoder-Decoder SLM, thereby improving the overall quality of the generated rationales.
3.4 Enhancing Task-Specific SLM via CoT Knowledge Distillation
@@
We conceptualize learning with rationales as a multi-task learning problem. Specifically, we train the model $f_{\omega}(x_{i})\to(y_{i},r_{i})$ to accomplish not just the prediction of task labels but also the generation of the corresponding rationales based on textual inputs. This multi-task training ensures that our model not only produces accurate predictions but also provides insightful justifications for its decisions. By doing so, we enhance the transparency and explainability of the model.
The multi-task learning objective for the Task-Specific SLM can be formulated as follows:
$\mathcal{L}_{2}=\mathcal{L}_{\text{Label}}+\mathcal{L}_{\text{Rationale}}$ (7)
@@
where $\ell_{\text{CE}}$ is the cross-entropy loss, $f_{\omega}(.)$ is the Task-Specific SLM model.
3.5 Privacy Analysis of FedCoT
The privacy-protection strategies in FedCoT implement a token-level Exponential Mechanism in feature space, adhering to the $\epsilon$-DP principles. This mechanism provides mathematically provable privacy guarantees at the token-level granularity, as extensively validated in privacy-preserving NLP research yue2021differential; chen-etal-2023-customized; tong2025inferdpt. Our experimental results further validate this approach: when privacy budget is low, the rationales generated from perturbed prompts show significantly lower similarity to those from original prompts, demonstrating the effectiveness of our privacy protection while acknowledging the inherent privacy-utility trade-off.
@@ 4.1 Setup
We have established a scenario to evaluate the performance of the FedCoT framework across a range of text generation tasks. This setup involves a client-server architecture, where the client holds two downstream SLMs: an Encoder-Decoder SLM, which specializes in encoder-decoder functionalities and a Task-Specific SLM, tailored for specific tasks. On the server-side, we host a LLM for more general and powerful text generation capabilities. Specifically, Table 1 outlines the detailed configurations of both the LLM and the SLMs. In our experimental setup, the Encoder-Decoder SLM and Task-Specific SLM are the identical architecture.
SLM
Setting LLM Encoder-Decoder Task-Specific
Setting 1 LLaMA3 70B Pythia-1.4B Pythia-1.4B
Setting 2 Qwen1.5-14B Qwen1.5-0.5B Qwen1.5-0.5B
Table 1: LLM and SLMs Setting of FedCoT.
Datasets and Evaluation Metrics. We conduct an evaluation of FedCoT on 4 QA datasets. Specifically, we include CommonsenseQA (CQA) talmor2019commonsenseqa, OpenBookQA (OBQA)  mihaylov2018can, BoolQ  clark2019boolq, ArcE clark2018think. For these datasets, we primarily use Accuracy as the evaluation metric. It’s worth noting that in our experiments, all methods undergo zero-shot evaluation except FewShot(1-shot), and we use the lm-evaluation-harness package eval-harness.
Baselines. Since we incorporate two distinct strategies in the prompt encoder and perturbed rationales decoder, we denote FedCoT method with the Exponential Mechanism Strategy as FedCoT-E and FedCoT method with the Adaptive Exponential Mechanism Strategy as FedCoT-A. We conduct a comparative analysis to evaluate the performance of our FedCoT framework, which comprises both FedCoT-E and FedCoT-A.
@@
  • Standalone, where the client fine-tunes its local model using its own private dataset;
  • Non-Private, where the client send its raw local prompt to server, get rationales from LLM and fine-tunes its local model like FedCoT, but without privacy-preserving.
4.2 Main Results
In this section, we undertake a comparative analysis of the task performance of FedCoT. We assess both the FedCoT-E and FedCoT-A methods against other baselines on Task-Specific SLM under the privacy budget $\epsilon=3$. Our experiments encompass two model configurations: Setting 1 (LLM: LLaMA3-70B, Encoder-Decoder SLM & Task-Specific SLM: Pythia-1.4B) and Setting 2 (LLM: Qwen1.5-14B, Encoder-Decoder SLM & Task-Specific SLM: Qwen1.5-0.5B).
The results, as presented in Table 2, clearly illustrate that both FedCoT-E and FedCoT-A exhibit significantly better performance when compared to FewShot-SLM and Standalone methods. Furthermore, FedCoT-A demonstrates notably superior performance compared to FedCoT-E. Specifically, take the model Setting 1 as an example, FedCoT-E surpasses the Standalone method by 4.3%, 3.2%, 7.1%, and 5.1% in the CQA, OBQA, BoolQ, and ArcE datasets, respectively. Meanwhile, FedCoT-A demonstrates even greater superiority, exceeding the Standalone method by 5.7%, 4.6%, 6.7%, and 6% across the same datasets.
Model Method CQA OBQA BoolQ ArcE
Setting 1
FewShot-LLM 70.29 80.66 90.08 82.69
FewShot-SLM 21.19 26.60 52.11 28.91
Standalone 42.43 38.73 73.07 40.33
Non-Private 49.22 46.07 80.61 48.01
FedCoT-E 46.70 41.93 80.02 45.42
FedCoT-A 48.10 43.30 79.77 46.34
Setting 2
@@ 4.2 Main Results
Non-Private 59.3 55.1 80.5 57.6
FedCoT-E 57.6 50.8 79 52.6
FedCoT-A 58.6 53.1 80.2 56.5
@@
Table 2: We compare the performance of Task-Specific SLM trained with FedCoT-E ($\epsilon=3$) and FedCoT-A ($\epsilon=3$) against the Task-Specific SLM trained using baseline methods. We consider two model settings: Setting 1 (LLM: LLaMA3-70B, Encoder-Decoder SLM & Task-Specific SLM: Pythia-1.4B) and Setting 2 (LLM: Qwen1.5-14B, Encoder-Decoder SLM & Task-Specific SLM: Qwen1.5-0.5B)
4.3 Performance Evaluation on various SLMs
In this section, we extend the evaluation of FedCoT’s effectiveness to encompass various client-side SLMs. These SLMs include LLaMA2-1.3B xia2024sheared, Qwen1.5-1.8B bai2023qwen, and OPT-1.3B zhang2022opt. We have chosen LLaMA3-70B dubey2024llama as LLM. Table 3 provides a clear illustration of how FedCoT(with $\epsilon$ = 3) consistently outperforms the Standalone method across various SLMs.
Dataset Method LLaMA2 Qwen1.5 OPT
CQA Standalone 61.5 57.8 56.42
FedCoT-E 63.03 60.30 57.55
FedCoT-A 64.27 62.21 60.18
OBQA Standalone 47.53 52.60 40.93
FedCoT-E 51.73 56.40 49.13
FedCoT-A 49.8 57.20 48.4
BoolQ Standalone 81.65 81.41 72.84
FedCoT-E 83.94 82.59 82.46
FedCoT-A 82.99 82.90 82.68
ArcE Standalone 40.33 55.58 45.92
FedCoT-E 54.11 61.07 49.67
FedCoT-A 54.66 62.43 50.69
Table 3: We compare the performance of Task-Specific SLMs, which have been trained with FedCoT-E($\epsilon=3$) and FedCoT-A($\epsilon=3$), against Standalone across various SLMs, including LLaMA2-1.3B, Qwen1.5-1.8B and OPT-1.3B.
4.4 Ablation Study
Influence of Privacy Budgets. We delve into the influence of privacy budgets on the performance of FedCoT. To ensure experimental consistency, we fix the model configuration to Setting 1 (as detailed in Table 1) for all subsequent ablation experiments. Table 4 presents an overview of FedCoT’s performance across a range of privacy budgets ($\epsilon=1,3,5,10$).
As the privacy budget $\epsilon$ increases, the performance of both FedCoT-E and FedCoT-A exhibits a notable uptick. Moreover, FedCoT-A consistently outperforms FedCoT-E under identical privacy budget conditions ($\epsilon$). When compared alongside Table 2, it becomes evident that with a privacy budget escalated to $\epsilon=10$, FedCoT-E surpasses the Standalone method by 5.6%, 6.1%, 6.3%, and 6.8% within the CQA, OBQA, BoolQ, and ArcE datasets, respectively. Similarly, FedCoT-A outperforms it by 4.3%, 7.1%, 6.8%, and 7%. Notably, across all evaluated datasets, at a privacy budget of $\epsilon=10$, FedCoT attains performance levels comparable to Non-Private approaches, underscoring its proficiency and adaptability in striking a balance between privacy and utility.
Method $\epsilon$ CQA OBQA BoolQ ArcE
FedCoT-E 1 45.63 42.13 78.91 44.84
3 46.70 41.93 80.02 45.42
5 46.50 43.35 80.17 46.70
10 48.03 44.87 79.37 47.14
FedCoT-A 1 47.31 43.20 79.63 46.65
3 48.10 43.30 79.77 46.34
5 47.96 44.20 79.91 48.08
10 47.74 45.81 79.86 47.30
Table 4: Comparison of the performance of Task-Specific SLM trained with FedCoT-E and FedCoT-A across different privacy budgets $\epsilon$.
Influence of Perturbed Rationales Decoding. We undertake an analysis to investigate the effects of perturbed rationales decoding on FedCoT when $\epsilon=3$. Table 5 offers a comparison of FedCoT’s performance, contrasting the results when perturbed rationales decoding is employed (FedCoT-E w/ and FedCoT-A w/) versus when it is not (FedCoT-E w/o and FedCoT-A w/o). Specifically, FedCoT-E w/ surpasses the FedCoT-E w/o by 2%, 1.3%, 1.5%, and 0.6% in the CQA, OBQA, BoolQ, and ArcE datasets, respectively. Meanwhile, FedCoT-A w/ demonstrates even greater superiority, exceeding the FedCoT-A w/o by 1.8%, 1.6%, 0.7%, and 3% across the same datasets. The findings unequivocally demonstrate that FedCoT exhibits superior performance when perturbed rationales decoding is utilized, as compared to when it is absent.
Decoding
Method Dataset w/ w/o
FedCoT-E CQA 46.70 44.79
OBQA 41.93 40.6
BoolQ 80.02 78.5
ArcE 45.42 44.78
FedCoT-A CQA 48.10 46.26
OBQA 43.30 41.7
BoolQ 79.77 79.06
ArcE 48.08 45.13
Table 5: Comparison of Task-Specific SLM Performance in FedCoT: With vs. Without perturbed rationales decoding.
@@ 4.4 Ablation Study
Perturbed Rationales vs Original Rationales. We focus on analyzing the quality of the perturbed rationales ($r^{p}$) generated from the perturbed prompt of LLM based on FedCoT-E and FedCoT-A methods and compare them with the rationales ($r$) generated from raw prompt of the LLM. To evaluate the similarity between $r^{p}$ and $r$, we use TokenRatio metric. A higher TokenRatio indicates a greater degree of similarity between the perturbed and original rationales.
@@
As shown in Table 6, with an increase in the privacy budget $\epsilon$ and a corresponding decrease in perturbation, both the TokenRatio of FedCoT-E and FedCoT-A have risen notably. Furthermore, in most of tasks, the TokenRatio of FedCoT-A is higher than that of FedCoT-E in the same level of privacy budget $\epsilon$. The experimental results confirm that the TokenRatio observed in the perturbed rationales produced by both FedCoT-E and FedCoT-A, positively correlate with the privacy budget $\epsilon$. This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the perturbed rationales become more similar to the original rationales.
Method $\epsilon$ CQA OBQA BoolQ ArcE
FedCoT-E 1 23.8 33 34.5 26.7
3 30.8 45.26 48.5 44.7
5 43.2 66.3 72.8 67.4
10 48.5 75.8 85.4 74.5
FedCoT-A 1 34.5 37.9 47.1 20.7
3 34.5 49.5 59.6 30
5 45.2 69.6 77.4 36.2
10 48.6 76.12 84.2 38.6
@@
Table 6: We conduct a comparative analysis to assess the perturbed rationales produced by FedCoT-E and FedCoT-A methods against the original rationales that are directly generated from the raw prompt of the LLM. Metric used: TokenRatio.
@@
Decoded Rationales vs Original Rationales. We delve into the quality analysis of the decoded rationales produced by the rationales decoder module based on FedCoT-E and FedCoT-A methods. We compare these decoded rationales against those generated directly from raw prompt of the LLM. We utilize the TokenRatio metric to assess their similarities.
@@
As shown in Table 7, in contrast to FewShot-SLM, it becomes apparent that the decoded rationales’ quality based on FedCoT-E and FedCoT-A methods isn’t solely reliant on the locally decoded SLM. The perturbed rationales crafted by the LLM indeed fulfill their intended purpose. When juxtaposed with Table 6, it’s clear that at comparable $\epsilon$ levels, the TokenRatio for the decoded rationales surpass those of the perturbed rationales in the FedCoT-E and FedCoT-A methods. This underscores the effectiveness of the rationales decoder module in the FedCoT-E and FedCoT-A methods.
Method $\epsilon$ CQA OBQA BoolQ ArcE
FewShot-SLM - 42.9 54.5 35.8 28.6
FedCoT-E 1 36 46.33 44.13 32.7
3 39 53.77 53.1 46
5 44.8 67.9 73.9 60.1
10 48.4 75.1 85.4 66.7
FedCoT-A 1 41.1 60.36 62.8 42.19
3 45.8 65.35 64.7 42.99
5 50 75.5 72.9 44.3
10 53.3 78.9 76.6 45.3
@@
Table 7: We conduct a comparative analysis to assess the decoded rationales produced by FedCoT-E and FedCoT-A methods against the original rationales that are directly generated from the raw prompt of the LLM. Metric used: TokenRatio.
Outperforming Standalone with 50% Data. We conduct an in-depth analysis to explore the influence of training data size on model performance. We compare the FedCoT method with the Standalone approach, varying the amount of training data used. Table 8 provides a clear illustration of how FedCoT(with $\epsilon=3$) achieves superior performance even with significantly fewer training samples compared to Standalone. More specifically, when trained on merely 50% of the complete CQA, OBQA, BoolQ, and ArcE datasets, both FedCoT-E and FedCoT-A either surpass or closely match the performance of Standalone method.
Dataset Method 25% 50% 75% 100%
CQA FedCoT-E 37.74 42.63 44.56 46.7
FedCoT-A 39.28 44.77 44.00 48.1
Standalone - - - 42.43
OBQA FedCoT-E 32.4 38.27 40.67 41.93
FedCoT-A 34.07 38.08 42.00 43.3
Standalone - - - 38.73
BoolQ FedCoT-E 69.96 72.26 77.67 80.02
FedCoT-A 69.61 73.73 77.82 79.77
Standalone - - - 73.07
ArcE FedCoT-E 37.79 41.42 42.22 45.42
FedCoT-A 37.64 41.86 45.28 46.34
Standalone - - - 40.33
Table 8: We compare the performance of Task-Specific SLM trained with FedCoT-E($\epsilon=3$) and FedCoT-A($\epsilon=3$) against Standalone, across a range of dataset sizes from 25% to 100%. The ’-’ indicates a method does not apply to the corresponding dataset sizes.